OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y less than or greater to x - 5 y greater than or equal to -x - 4

OpenStudy (johnweldon1993):

Provide us with the points...

OpenStudy (anonymous):

(-5, 2) (5, -2) (-5, -2) (5, 2)

OpenStudy (johnweldon1993):

Alright...so you have some points...in the form (x , y) and you have 2 equations y <= x - 5 y >= -x - 4 So what you would do...is plug in those numbers you have (x , y) in for both 'x' and 'y' in your equations...and see which on works out...here I'll do the first choice for an example...

OpenStudy (anonymous):

Alright

OpenStudy (johnweldon1993):

the point (-5 , 2) the equations y <= x - 5 y >= -x - 4 after plugging in our point we have...we have the 2 equations that look like 2 <= -5 - 5 2 >= -(-5) - 4 lets see if they work 2 <= -10 *is 2 less than or equal to -10? no we already know this choice will not work 2 >= 1 *is 2 greater than or equal to 1? yes...but look above...the other one doesn't work so this choice is NOT the answer

OpenStudy (johnweldon1993):

And you would do this same process with the other choices you have...

OpenStudy (anonymous):

Okay great. Thanks a million hun